the length of a rectangle is 3 feet more than twice the width. the area is 90 square feet. what are the dimensions of the rectangle?

Answer by ReadingBoosters(3246)

a = lw
Equations
l = 3 + 2w
90 = lw
Substitute the first into the second equation
90 = (3+2w)w
90 = 3w + 2w^2
2w^2 + 3w - 90 = 0
Multiples of 90: 30,3; 45,2; 18,5; 6,15
Looking for combination where 2 x one of the pair will subtract from the other to yield 3
(2w + 15)(w - 6) Mixed signs due to - 90
2w + 15 = 0; w = -15/2 NO
w - 6 = 0; w = 6
width = 6
length = 3 + 2(6) = 15